# LongCat-Flash-Omni Technical Report

### v1

@@ 1 Introduction
Humans are inherently omni-modal beings, capable of efficiently perceiving and integrating diverse forms of information, including visual and auditory inputs, to accomplish a wide range of challenging tasks. The seamless combination and transmission of multiple modalities significantly enhance the effectiveness and efficiency of human communication and interaction. In pursuit of Artificial General Intelligence (AGI), the field of large language models (LLMs) is now rapidly evolving toward the integration of richer multimodal capabilities and more efficient human-AI interaction. Recent pioneers such as Gemini-2.5 (Comanici et al., 2025) and GPT-4o (OpenAI, 2024) have integrated text, audio, image, and video processing within a single model, enabling efficient audio-visual interaction. Following these advances, research on omni-modal models has attracted broad attention, with many subsequent efforts proposed in the community (Xu et al., 2025; Guo et al., 2025; Li et al., 2025; Liu et al., 2025; Fu et al., 2025).
@@
To address the third challenge of achieving low-latency audio-visual interaction in a large-scale model, we dedicate substantial effort to the design of all modules in LongCat-Flash-Omni. We adopt the ScMoE architecture with zero-computation experts from LongCat-Flash (Meituan, 2025a, b) as the LLM backbone. To handle streaming input, we employ efficient audio and video encoders for feature extraction and introduce a synchronized chunk-wise interleaving strategy for real-time processing. For efficient audio reconstruction, we adopted a multi-codebook audio detokenization scheme with a coarser temporal resolution. This approach significantly improves decoding efficiency while preserving reconstruction quality. Furthermore, we designed an efficient streaming pipeline for model serving to minimize end-to-end server-side latency. As a result, despite having up to 560B parameters (with 27B activated), LongCat-Flash-Omni achieves millisecond-level response latency in real-time interactive scenarios.
@@ 2 Architecture
As illustrated in Figure 2, LongCat-Flash-Omni is a fully end-to-end omni-modal model. It can receive various modalities as input, including text, audio, image, video, and their any arbitrary combination, and is able to directly generates speech tokens from the LLM backbone. LongCat-Flash-Omni employs a vision encoder and an audio encoder as multimodal perceivers. An LLM processes the multimodal inputs and generates text and audio tokens. An audio decoder reconstructs the waveform from the speech tokens generated by the LLM, enabling natural speech interaction. All modules are carefully designed to support efficient streaming inference. The audio encoder, vision encoder, and audio decoder are lightweight components, each with approximately 600M parameters. The large-scale LLM backbone uses the novel and efficient architectural design proposed in the LongCat LLM family (Meituan, 2025a, b).
@@ 2.1 Vision Encoder
The vision encoder serves as a critical component in multimodal language models. To effectively encode visual inputs such as images and videos, LongCat-Flash-Omni incorporates a well-designed Vision Transformer (ViT), referred to as LongCat-ViT (Qiao et al., 2025). LongCat-ViT achieves high performance across multimodal tasks, natively supports inputs with various resolutions and aspect ratios, while providing unified encoding capabilities for both image and video data.
@@
Native Resolution Encoding Conventional ViT models widely adopted in the community (e.g., CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023)) typically resize input images to a fixed resolution, which can result in substantial information loss, particularly for images with extreme aspect ratios or high native resolutions. To mitigate this limitation, LongCat-ViT encodes visual inputs at their native resolutions, circumventing the limitations of fixed-resolution ViT models. This preserves the spatial and contextual information inherent in visual data, thereby enhancing the model’s capability to comprehend and reason over complex visual inputs. For each image or video frame, if the number of patches falls within a predefined range (576-5832 during training), only minimal resizing is applied to ensure both dimensions are divisible by 112. Otherwise, the image is rescaled to fit within this range while preserving its aspect ratio.
@@ 2.2 Audio Tokenizer, Encoder, and Decoder
Audio Tokenizer and Decoder We adopt LongCat-Audio-Codec (Zhao et al., 2025a) as our audio tokenizer and decoder, owing to its robust semantic modeling, flexible acoustic feature extraction, and low-latency streaming synthesis capabilities. The tokenizer discretizes audio waveforms into four codebooks at a frame rate of 16.67 Hz, with one codebook representing semantic information and the other three capturing acoustic details. To achieve low-latency inference in real-time interaction scenarios, unlike conventional waveform reconstruction methods that rely on diffusion- or flow-matching-based code2mel models followed by vocoders, we directly employ the decoder from LongCat-Audio-Codec as the audio decoder to reconstruct waveform from tokens. It supports streaming audio decoding with a look-ahead of only three frames. As illustrated in Figure 4, the audio decoder consists of LSTM layers, convolutional blocks, and causal transposed convolution layers, and is trained under a generative adversarial network (GAN) framework.
@@
Audio Encoder To optimize response latency and accommodate speech inputs of arbitrary duration, the audio encoder is designed using a streaming architecture. As illustrated in Figure 4, the audio encoder takes 80-dimensional Fbank features as input. The architecture incorporates a Pre-FFN module that reduces the audio sequence length by a factor of eight through frame splicing downsampling technique, with each frame representing an 80ms time window. The core process is performed by a streaming encoder that maintains a Transformer-like structure while incorporating several key modifications: (1) a Pre-Norm configuration for enhanced training stability, and (2) replacing standard self-attention modules with FSMN layers (Zhang et al., 2018) to enable efficient feature processing within constrained context windows. To balance latency and performance, we implement a hybrid approach where only the final six layers incorporate a one-frame look-ahead mechanism, while maintaining strict causality in the preceding layers. The architecture concludes with a post-FFN module for additional feature refinement. The audio encoder is trained under supervised learning using speech recognition data using the CTC loss (Graves et al., 2006).
@@ 2.3 LLM Backbone
LongCat-Flash-Omni is built upon LongCat-Flash (Meituan, 2025a), a 560-billion-parameter Mixture-of-Experts (MoE) language model. LongCat-Flash adopts Multi-head Latent Attention (MLA) (Liu et al., 2024a), shortcut-connected MoE (Cai et al., 2024) and zero-computation experts, performing variable computation per token by activating 18.6B-31.3B parameters (27B on average), thereby unifying efficiency, performance and sparsity. These characteristics are retained and extended to multimodal understanding and audio-visual interaction by LongCat-Flash-Omni.
@@ 3.1.1 Audio Data
Speech-Text Interleaved Data Our data consists of tens of millions of hours of audio with rich diversity. We employ the following pipeline to extract high-quality speech audios with consistent topics: First, we use a VAD system to split a long audio into speaking segments and remove non-speaking regions. Next, we use two proprietary ASR models for cross-validation, filtering out segments with significant discrepancies on transcription results. A multilingual speech aligner (Pratap et al., 2024) is then applied for forced alignment, yielding precise transcription timestamps. We further refine the dataset by computing the ratio between speech duration and text length for each segment, and discard the segments whose ratios fall outside the $0.5$-th to $99.5$-th percentile range. Finally, we merge adjacent speaking segments separated by silences shorter than 10 seconds to form the training samples.
@@ 3.1.2 Generic Image-Text Data
To broaden semantic coverage and avoid overfitting to dominant patterns, we significantly enhance data diversity and balance. We cluster image-text pairs based on joint image and text embeddings and resample from each cluster to ensure representation of long-tail content, while simultaneously removing low-quality samples frequently concentrated in specific clusters. Experimental results confirm that preserving diversity at this stage is more beneficial than overly strict quality filtering. Furthermore, to address the pronounced long-tail distribution inherent in multimodal datasets, we adopt a concept-based resampling strategy inspired by MetaCLIP (Xu et al., 2023). By expanding the vocabulary with a 200K-scale Chinese lexicon and resampling for broader concept coverage, we achieve a more balanced distribution across semantic categories.
Interleaved Image-Text Data Interleaved image-text data provides broader visual-textual coverage and improves few-shot performance, but it is often noisy and has uneven quality. To address this, we construct a high-quality dataset through a two-stage pipeline of filtering and diversity sampling from open-source data. In the filtering stage, we remove samples with noisy tokens, sensitive content, and overly complex samples, and discard corrupted or low-resolution images. We then improve image-text alignment by pairing each image with its most relevant text segment by SigLIP similarity scores, and discard pairs with low semantic alignment. To sample a diverse and evenly distributed subset, we apply density-based pruning and semantic clustering similar to (Abbas et al., 2024). Samples are scored by knowledge content, quality, and similarity, then evenly sampled based on the cluster they belong to and their distances to the cluster centroid. This process reduces the raw dataset by approximately 74%, while maintaining diversity and quality for multimodal pretraining.
@@ 3.1.3 OCR, Grounding and GUI Data
Grounding Data To enhance the model’s grounding capabilities, we adopt several widely used open-source object detection datasets (Krishna et al., 2017; Mao et al., 2016; Lin et al., 2014; Shao et al., 2019). We perform data quality validation to filter out incorrect and redundant samples. Using these high-quality open-source datasets, we construct two types of question-answering data: localization data and region-captioning data. We uniformly adopt relative coordinates, which are normalized to the range of 0-1000, and process the data into JSON format. We also incorporate PixMo dataset (Deitke et al., 2024) to enhance the model’s counting ability.
GUI Data Graphical User Interface (GUI) data contains rich information on visual understanding and task planning, enabling automated interactions on both mobile and desktop platforms. To this end, we utilize diverse types of GUI-related data to enhance the model’s perception, grounding, and planning capabilities (Zeng et al., 2025). For GUI perception, we utilize a large number of image-text pairs from screenshots of various PC and mobile applications, covering tasks such as OCR, VQA, and captioning with GUI images. For GUI grounding, we construct instruction-answer pairs based on visual screenshots from diverse platforms and devices. For each screenshot we create multiple pairs, accounting for different UI elements and interaction possibilities. For GUI planning, we collect a diverse set of navigational paths with rich context from both mobile and desktop. The planning data comprises key components such as screenshot observations, summarizations, reasoning traces, and corresponding actions. To enhance the model’s reasoning capability, we integrate multiple inference settings that involve different combinations of action, summarization, and reasoning outputs. Furthermore, to capture the underlying logic of GUI dynamics, we incorporate contextual scenarios that enable the model to predict changes across different screenshots.
@@ 3.1.6 Video Data
High-quality video caption data Our video captioning process begins by dividing videos into coherent scenes using a scene detection algorithm similar to Koala (Wang et al., 2025a). To ensure a wide range of content, we then cluster and sample these scenes. Finally, we employ a proprietary model to generate detailed captions that describe the sequence of events and the overall context of the video.
@@
Action recognition video QA data To strengthen the model’s action recognition capability, we transform several publicly available video recognition datasets (Carreira et al., 2019; Goyal et al., 2017) into multiple-choice and open-ended video QA pairs.
@@ 3.2.1 Stage-0 Text Pre-Training
To establish a robust text foundation, the text pre-training stage follows the same procedure as the initial phase of LongCat-Flash (Meituan, 2025a). The model is trained on approximately 16 trillion tokens drawn from a high-quality and diverse text corpus, where a constant learning rate is applied. Throughout training, the proportion of high-quality reasoning data (e.g., STEM and code) is progressively increased to strengthen the model’s reasoning and programming capabilities, thereby providing a solid foundation for subsequent multimodal learning.
@@ 3.2.4 Stage-3 Multimodal Annealing Training
We adopt a perplexity (PPL)-gap-based signal to automatically guide data sampling allocation. The corpus is first segmented into distinct subsets according to semantics and tasks. During training, we monitor PPL convergence for each subset: if a subset’s convergence lags behind the expected reference level, its sampling weight will be dynamically increased. For each subset, we build a corresponding validation set and use an off-the-shelf vision-language model to compute its per-sample PPL as the expected reference level (Xie et al., 2023; Michel et al., 2022). To prevent high-value samples from being diluted by aggregate statistics, samples for which the PPL convergence is consistent with downstream performance will be isolated and relabeled into independent data subsets. With the original data partitions largely preserved while the high-value samples organized more finely, this approach significantly enhances data efficiency and downstream task performance.
@@ 3.2.5 Stage-4 Context-Length Extension Training
To augment the model’s proficiency in capturing extended sequential relationships across various modalities, we gradually extend the context length to 32K and then 128K tokens, enabling the model to process and reason over more complex tasks such as long-term memory modeling and multi-turn real-time interaction. We scale the context length from 8K to 32K tokens using 100B training tokens and adjust RoPE’s base frequency (Su et al., 2024) from 1M to 5M to preserve positional encoding quality. The context length is then further expanded to 128K tokens with an additional 20B tokens of training, requiring a proportional increase of the RoPE’s base frequency to 10M to maintain stable attention across the extended sequence length.
@@
For textual and speech modalities, we utilize the foundation language model’s high-quality text corpus (Meituan, 2025a), alongside carefully curated speech data from our multilingual audio repository (over 10 million hours of processed Chinese/English recordings). Notably, we maintain the same 2:1:1 text-to-vision-to-speech ratio as established in the pre-training phase, ensuring modality balance during context extension. This comprehensive approach ensures consistent performance across all modalities while scaling to extended contexts.

### v2

@@ 1 Introduction
Humans are inherently omni-modal beings, capable of efficiently perceiving and integrating diverse forms of information, including visual and auditory inputs, to accomplish a wide range of challenging tasks. The seamless combination and transmission of multiple modalities significantly enhance the effectiveness and efficiency of human communication and interaction. In pursuit of Artificial General Intelligence (AGI), the field of large language models (LLMs) is now rapidly evolving toward the integration of richer multimodal capabilities and more efficient human-AI interaction. Recent pioneers such as Gemini-2.5 (comanici2025gemini) and GPT-4o (gpt4o) have integrated text, audio, image, and video processing within a single model, enabling efficient audio-visual interaction. Following these advances, research on omni-modal models has attracted broad attention, with many subsequent efforts proposed in the community (xu2025qwen3; guo2025m2; li2025baichuan; liu2025ola; fu2025vita).
@@
To address the third challenge of achieving low-latency audio-visual interaction in a large-scale model, we dedicate substantial effort to the design of all modules in LongCat-Flash-Omni. We adopt the ScMoE architecture with zero-computation experts from LongCat-Flash (meituan2025longcat_flash_chat; meituan2025longcat_flash_chat_thinking) as the LLM backbone. To handle streaming input, we employ efficient audio and video encoders for feature extraction and introduce a synchronized chunk-wise interleaving strategy for real-time processing. For efficient audio reconstruction, we adopted a multi-codebook audio detokenization scheme with a coarser temporal resolution. This approach significantly improves decoding efficiency while preserving reconstruction quality. Furthermore, we designed an efficient streaming pipeline for model serving to minimize end-to-end server-side latency. As a result, despite having up to 560B parameters (with 27B activated), LongCat-Flash-Omni achieves millisecond-level response latency in real-time interactive scenarios.
@@ 2 Architecture
As illustrated in Figure 2, LongCat-Flash-Omni is a fully end-to-end omni-modal model. It can receive various modalities as input, including text, audio, image, video, and their any arbitrary combination, and is able to directly generates speech tokens from the LLM backbone. LongCat-Flash-Omni employs a vision encoder and an audio encoder as multimodal perceivers. An LLM processes the multimodal inputs and generates text and audio tokens. An audio decoder reconstructs the waveform from the speech tokens generated by the LLM, enabling natural speech interaction. All modules are carefully designed to support efficient streaming inference. The audio encoder, vision encoder, and audio decoder are lightweight components, each with approximately 600M parameters. The large-scale LLM backbone uses the novel and efficient architectural design proposed in the LongCat LLM family (meituan2025longcat_flash_chat; meituan2025longcat_flash_chat_thinking).
@@ 2.1 Vision Encoder
The vision encoder serves as a critical component in multimodal language models. To effectively encode visual inputs such as images and videos, LongCat-Flash-Omni incorporates a well-designed Vision Transformer (ViT), referred to as LongCat-ViT (qiao2025univitar). LongCat-ViT achieves high performance across multimodal tasks, natively supports inputs with various resolutions and aspect ratios, while providing unified encoding capabilities for both image and video data.
@@
Native Resolution Encoding Conventional ViT models widely adopted in the community (e.g., CLIP (radford2021clip), SigLIP (zhai2023sigmoid)) typically resize input images to a fixed resolution, which can result in substantial information loss, particularly for images with extreme aspect ratios or high native resolutions. To mitigate this limitation, LongCat-ViT encodes visual inputs at their native resolutions, circumventing the limitations of fixed-resolution ViT models. This preserves the spatial and contextual information inherent in visual data, thereby enhancing the model’s capability to comprehend and reason over complex visual inputs. For each image or video frame, if the number of patches falls within a predefined range (576-5832 during training), only minimal resizing is applied to ensure both dimensions are divisible by 112. Otherwise, the image is rescaled to fit within this range while preserving its aspect ratio.
@@ 2.2 Audio Tokenizer, Encoder, and Decoder
Audio Tokenizer and Decoder We adopt LongCat-Audio-Codec (longcat-codec2025zhao) as our audio tokenizer and decoder, owing to its robust semantic modeling, flexible acoustic feature extraction, and low-latency streaming synthesis capabilities. The tokenizer discretizes audio waveforms into four codebooks at a frame rate of 16.67 Hz, with one codebook representing semantic information and the other three capturing acoustic details. To achieve low-latency inference in real-time interaction scenarios, unlike conventional waveform reconstruction methods that rely on diffusion- or flow-matching-based code2mel models followed by vocoders, we directly employ the decoder from LongCat-Audio-Codec as the audio decoder to reconstruct waveform from tokens. It supports streaming audio decoding with a look-ahead of only three frames. As illustrated in Figure 4, the audio decoder consists of LSTM layers, convolutional blocks, and causal transposed convolution layers, and is trained under a generative adversarial network (GAN) framework.
@@
Audio Encoder To optimize response latency and accommodate speech inputs of arbitrary duration, the audio encoder is designed using a streaming architecture. As illustrated in Figure 4, the audio encoder takes 80-dimensional Fbank features as input. The architecture incorporates a Pre-FFN module that reduces the audio sequence length by a factor of eight through frame splicing downsampling technique, with each frame representing an 80ms time window. The core process is performed by a streaming encoder that maintains a Transformer-like structure while incorporating several key modifications: (1) a Pre-Norm configuration for enhanced training stability, and (2) replacing standard self-attention modules with FSMN layers (zhang2018deep) to enable efficient feature processing within constrained context windows. To balance latency and performance, we implement a hybrid approach where only the final six layers incorporate a one-frame look-ahead mechanism, while maintaining strict causality in the preceding layers. The architecture concludes with a post-FFN module for additional feature refinement. The audio encoder is trained under supervised learning using speech recognition data using the CTC loss (graves2006connectionist).
@@ 2.3 LLM Backbone
LongCat-Flash-Omni is built upon LongCat-Flash (meituan2025longcat_flash_chat), a 560-billion-parameter Mixture-of-Experts (MoE) language model. LongCat-Flash adopts Multi-head Latent Attention (MLA) (liu2024deepseek), shortcut-connected MoE (cai2024shortcut) and zero-computation experts, performing variable computation per token by activating 18.6B-31.3B parameters (27B on average), thereby unifying efficiency, performance and sparsity. These characteristics are retained and extended to multimodal understanding and audio-visual interaction by LongCat-Flash-Omni.
@@ 3.1.1 Audio Data
Speech-Text Interleaved Data Our data consists of tens of millions of hours of audio with rich diversity. We employ the following pipeline to extract high-quality speech audios with consistent topics: First, we use a VAD system to split a long audio into speaking segments and remove non-speaking regions. Next, we use two proprietary ASR models for cross-validation, filtering out segments with significant discrepancies on transcription results. A multilingual speech aligner (pratap2024scaling) is then applied for forced alignment, yielding precise transcription timestamps. We further refine the dataset by computing the ratio between speech duration and text length for each segment, and discard the segments whose ratios fall outside the $0.5$-th to $99.5$-th percentile range. Finally, we merge adjacent speaking segments separated by silences shorter than 10 seconds to form the training samples.
@@ 3.1.2 Generic Image-Text Data
To broaden semantic coverage and avoid overfitting to dominant patterns, we significantly enhance data diversity and balance. We cluster image-text pairs based on joint image and text embeddings and resample from each cluster to ensure representation of long-tail content, while simultaneously removing low-quality samples frequently concentrated in specific clusters. Experimental results confirm that preserving diversity at this stage is more beneficial than overly strict quality filtering. Furthermore, to address the pronounced long-tail distribution inherent in multimodal datasets, we adopt a concept-based resampling strategy inspired by MetaCLIP (xu2023demystifying). By expanding the vocabulary with a 200K-scale Chinese lexicon and resampling for broader concept coverage, we achieve a more balanced distribution across semantic categories.
Interleaved Image-Text Data Interleaved image-text data provides broader visual-textual coverage and improves few-shot performance, but it is often noisy and has uneven quality. To address this, we construct a high-quality dataset through a two-stage pipeline of filtering and diversity sampling from open-source data. In the filtering stage, we remove samples with noisy tokens, sensitive content, and overly complex samples, and discard corrupted or low-resolution images. We then improve image-text alignment by pairing each image with its most relevant text segment by SigLIP similarity scores, and discard pairs with low semantic alignment. To sample a diverse and evenly distributed subset, we apply density-based pruning and semantic clustering similar to (abbas2024effective). Samples are scored by knowledge content, quality, and similarity, then evenly sampled based on the cluster they belong to and their distances to the cluster centroid. This process reduces the raw dataset by approximately 74%, while maintaining diversity and quality for multimodal pretraining.
@@ 3.1.3 OCR, Grounding and GUI Data
Grounding Data To enhance the model’s grounding capabilities, we adopt several widely used open-source object detection datasets (krishna2017visualgenome; mao2016refcocoplusg; lin2014microsoftcoco; shao2019objects365). We perform data quality validation to filter out incorrect and redundant samples. Using these high-quality open-source datasets, we construct two types of question-answering data: localization data and region-captioning data. We uniformly adopt relative coordinates, which are normalized to the range of 0-1000, and process the data into JSON format. We also incorporate PixMo dataset (deitke2024molmo) to enhance the model’s counting ability.
GUI Data Graphical User Interface (GUI) data contains rich information on visual understanding and task planning, enabling automated interactions on both mobile and desktop platforms. To this end, we utilize diverse types of GUI-related data to enhance the model’s perception, grounding, and planning capabilities (zeng2025uitron). For GUI perception, we utilize a large number of image-text pairs from screenshots of various PC and mobile applications, covering tasks such as OCR, VQA, and captioning with GUI images. For GUI grounding, we construct instruction-answer pairs based on visual screenshots from diverse platforms and devices. For each screenshot we create multiple pairs, accounting for different UI elements and interaction possibilities. For GUI planning, we collect a diverse set of navigational paths with rich context from both mobile and desktop. The planning data comprises key components such as screenshot observations, summarizations, reasoning traces, and corresponding actions. To enhance the model’s reasoning capability, we integrate multiple inference settings that involve different combinations of action, summarization, and reasoning outputs. Furthermore, to capture the underlying logic of GUI dynamics, we incorporate contextual scenarios that enable the model to predict changes across different screenshots.
@@ 3.1.6 Video Data
High-quality video caption data Our video captioning process begins by dividing videos into coherent scenes using a scene detection algorithm similar to Koala (wang2025koala). To ensure a wide range of content, we then cluster and sample these scenes. Finally, we employ a proprietary model to generate detailed captions that describe the sequence of events and the overall context of the video.
@@
Action recognition video QA data To strengthen the model’s action recognition capability, we transform several publicly available video recognition datasets (carreira2019short; Goyal_2017_ICCV) into multiple-choice and open-ended video QA pairs.
@@ 3.2.1 Stage-0 Text Pre-Training
To establish a robust text foundation, the text pre-training stage follows the same procedure as the initial phase of LongCat-Flash (meituan2025longcat_flash_chat). The model is trained on approximately 16 trillion tokens drawn from a high-quality and diverse text corpus, where a constant learning rate is applied. Throughout training, the proportion of high-quality reasoning data (e.g., STEM and code) is progressively increased to strengthen the model’s reasoning and programming capabilities, thereby providing a solid foundation for subsequent multimodal learning.
@@ 3.2.4 Stage-3 Multimodal Annealing Training
We adopt a perplexity (PPL)-gap-based signal to automatically guide data sampling allocation. The corpus is first segmented into distinct subsets according to semantics and tasks. During training, we monitor PPL convergence for each subset: if a subset’s convergence lags behind the expected reference level, its sampling weight will be dynamically increased. For each subset, we build a corresponding validation set and use an off-the-shelf vision-language model to compute its per-sample PPL as the expected reference level (DBLP:conf/nips/Xie0DDLLLL0Y23; DBLP:conf/iclr/MichelHN22). To prevent high-value samples from being diluted by aggregate statistics, samples for which the PPL convergence is consistent with downstream performance will be isolated and relabeled into independent data subsets. With the original data partitions largely preserved while the high-value samples organized more finely, this approach significantly enhances data efficiency and downstream task performance.
@@ 3.2.5 Stage-4 Context-Length Extension Training
To augment the model’s proficiency in capturing extended sequential relationships across various modalities, we gradually extend the context length to 32K and then 128K tokens, enabling the model to process and reason over more complex tasks such as long-term memory modeling and multi-turn real-time interaction. We scale the context length from 8K to 32K tokens using 100B training tokens and adjust RoPE’s base frequency (su2024roformer) from 1M to 5M to preserve positional encoding quality. The context length is then further expanded to 128K tokens with an additional 20B tokens of training, requiring a proportional increase of the RoPE’s base frequency to 10M to maintain stable attention across the extended sequence length.
@@
For textual and speech modalities, we utilize the foundation language model’s high-quality text corpus (meituan2025longcat_flash_chat), alongside carefully curated speech data from our multilingual audio repository (over 10 million hours of processed Chinese/English recordings). Notably, we maintain the same 2:1:1 text-to-vision-to-speech ratio as established in the pre-training phase, ensuring modality balance during context extension. This comprehensive approach ensures consistent performance across all modalities while scaling to extended contexts.